OpenStudy (anonymous):

write the expression in terms of sin x and cos x only. sin(2x)+cos(3x)

OpenStudy (anonymous):

cos3x = cos (2x+ x) =cos2x*cosx - sin2x*sinx =(2*cos^2 (x) - 1) *cosx - 2sinx*cosx*sinx { since cos2x = 2cos^2 (X) -1 , sin2x = 2*sinx*cosx } =2 cos^3 (x) - cos x - 2 cosx *sin^2 (x) = 2 cos^3 (x) - cos x (1 + 2 sin^2 (x) ) = 2 cos^3 (x) - cos x (1 + 2 ( 1 - cos^2 (x) )) = 2 cos^3 (x) - cos x ( 1+ 2 - 2*cos^2(x) ) =2 cos^3 (x) - cos x (3 -2*cos^2(x) ) =2 cos^3 (x) - 3cos x + 2*cos^3(x) =4 cos^3 (x) - 3cos x

OpenStudy (anonymous):

sin(2x+x) = cos(2x)sin(x) + sin(2x)cos(x) Then substitute the double angle formulae for the 2xs = (cos^2(x) - sin^2(x))sin(x) + 2cos(x)sin(x)cos(x) // factor out the sinx = sin(x) (cos^2(x)-sin^2(x) + 2 cos^2(x)) =sin(x) (3cos^2(x) - sin^2(x))

OpenStudy (anonymous):

but for the second one it was sin(2x) not sin(3x) and using the identity the sin(2x) can become 2sin(x)cos(x) when would make the problem turn into 2sin(x)cos(x)+4 cos^3 (x) - 3cos x is that the final answer? or- I'm actually kinda confused on what the question is asking, does the answer need to using just sin or just cos, or can it use both? I copied the question word for word from the book onto here

OpenStudy (anonymous):

2sin(x)cos(x)+4 cos^3 (x) - 3cos x is the final answer